The table shows the demand and supply schedules for a good before and after the imposition of a tax.

price quantity demanded quantity supplied before tax quantity supplied after tax
20 340 440 380
19 340 430 340
18 340 410 290
17 340 380 230
16 340 340 160
15 340 290 80
14 340 230 0
What was the amount of the tax revenue raised for the government?
A $1020 B $1360 C $5440 D $6460
1
Expert's answer
2016-05-30T09:42:03-0400
Before tax the equilibrium quantity was Qd = Qs = 340 units, P = $16. After tax Qd = Qs = 340 too and P = $19, so the amount of tax is 19 - 16 = $3 per unit. As Q = 340 units, then the amount of the tax revenue raised for the government is Tax Revenue = t*Q = 3*340 = $1020.
So, the correct answer is A $1020.